OpenStudy (anonymous):

If the probability density of a random vribale is given by : f(x)= x for 0

OpenStudy (anonymous):

My issue isn't how to integrate between the two, but the process in finding the answer. Do I have to integrate each function first, plug in the limits from x = 0 to 1 (for function 1), x=1 to 2 (for function 2) and then integrate that answer for each function and plug in the two probabilities for both functions?

OpenStudy (anonymous):

^^Oh and that word should say "variable" in my first post.

OpenStudy (anonymous):

If I'm right then would I have to do the following: \[f(x)=\int\limits_{0}^{1}x dx= 1/2\] \[f(x)=\int\limits_{.2}^{.8}1/2dx=.8/2-.2/2=.3\]

OpenStudy (zarkon):

.3 is good

OpenStudy (zarkon):

though your notation and work is poor

OpenStudy (zarkon):

\[\int\limits_{.2}^{.8}xdx=.3\]

OpenStudy (zarkon):

Just to see the notation... \[P(.2<X<.8)=\int\limits_{.2}^{.8}f(x)dx=\int\limits_{.2}^{.8}xdx=.3\]

OpenStudy (anonymous):

So I didn't have to find the definite integral of: \[\int\limits_{0}^{1}xdx\]?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

if \(X\) has density \(f(x)\) then \[P(a<X<b)=\int\limits_{a}^{b}f(x)dx\]

OpenStudy (anonymous):

Oh ok. Thanks, I think I can do the rest from here.